# Social perception of faces in a vision-language model

### v1

@@ B.1 Adapted SC-WEAT
Tab. S.3 shows the WEAT scores for CausalFace and the two observational datasets. The most striking similarities in means are observed when comparing Asians to Blacks and Whites to Blacks. In the comparison Male versus Female, CausalFace and FairFace demonstrate closely matched scores, while UTKFace records a lower WEAT value.
B.2 Markedness
@@
B.3 Mean cosine similarities
@@
B.4 Skew@k, MaxSkew@k, NDKL
We present three additional bias metrics, following [21].
@@ B.4 Skew@k, MaxSkew@k, NDKL
The most left block of the bottom subtable in Tab. S.3 depicts Skew@k. Asians predominantly receive positive values, indicative of a favorable social perception relative to other races. The values for Blacks, however, are distinctly negative. In the case of Whites, CausalFace consistently shows negative social perception, while FairFace and UTKFace mostly hover around zero. In contrast, Skew@k for men and women hovers near zero, implying no significant bias towards either gender.
@@
Appendix C Additional results
C.1 Facial expression, lighting, and pose confound social perception
@@
Fig. S.1 illustrates the impact of three non-protected characteristics—facial expression, lighting, and pose—on various social perception dimensions. The cosine similarity of a neutral text prompt is depicted by the gray horizontal dashed line, serving as a reference. Deviations from this reference, caused by adding a valence attribute to text prompts, are represented as markers connected by gray lines. To support the visual interpretation, we calculated Pearson correlation coefficients $r$ (Tab. S.4).
As we are interested in variables that provoke differences in social perceptions, we focus on facial expressions in the main paper. In the following, we interpret lighting and pose in more detail.
@@
The middle panel of Fig. S.1 illustrates the effect of lighting variations on social perception, revealing a consistent pattern across all dimensions. These variations, represented on the x-axis, correspond to different angles of light incidence, visualized in Fig. S.11. It is important to note that these variations should not be understood linearly, despite their representation on a linear scale for comparison purposes across the three non-protected characteristics. Thus, the slope of the curves in this plot is not a focal point of our analysis. For all social dimensions, the highest scores are observed when the face is evenly lit (lighting = -1). Other configurations with high scores include lighting from above (lighting = 2 and 3). These findings are consistent with human subject studies [72, 73]. In contrast, the lowest scores are observed when the face is lit from below (lighting = 5 and 6), which tends to be perceived as unnatural by humans.
@@
The right panel of Fig. S.1 shows variations of $\Delta$ cosine similarity in pose. We observe a parallel trend in all social perception dimensions. The parallel trend is mirrored by a positive $r$ for all dimensions (average: $r$=0.79). The curves are mirrored around a pose value of zero, representing a frontal stance. As a face tilts more to the left (pose $<$ 0) or to the right (pose $>$ 0), the curves approach closer to the baseline. This effect is easier to interpret when considering that the raw (not the $\Delta$) cosine similarities instead. The raw values are the highest for frontal poses and decrease the more an image is tilted to the left or right (shaped like an upward-facing triangle). In other words, for side-pose photos, the image embedding is more distant from the “person” mode in the embedding space. This makes the interpretation of changes induced by adding social adjectives to the text prompt challenging. This observation suggests that uncontrolled pose variations in the observational, wild-collected datasets are a significant source of noise that makes analyzing other effects difficult and could explain a significant share of the noise present in wild-collected datasets (Appx. E). Future research could address this observation in more detail. Additionally, psychological literature indicates that head orientation (pose) can influence perceived emotions [74]. This would suggest that we should find a different impact on positive and negative attributes, at least for some of the varied head positions. However, in [74], the postures affecting emotions were upward and downward head postures, while in CasualFace, we have left-right orientation variation, which may explain the difference in our observations. Additionally, we observe that negative valence is consistently ranked lower than positive valence, unaffected by head position. This suggests that another factor, currently unmeasured and uncontrolled, may affect this aspect.
@@
Figure S.1: Social perception as a function of Smiling, Lighting, and Pose. Curves for W (Warmth), C (Competence), A+, A- (Agency), BP, BC (Belief), C+, C- (Communion). Visualizations of the x-axis values are shown in Figs. S.9, S.10 and S.11. The cosine similarity between an image and a neutral text prompt is used as the reference and is depicted by the horizontal gray dashed line at 0. Markers are obtained by averaging across races and genders. Please see Figs. 3,5 for intersectional plots.
@@
C.2 Ageing influences social perception across intersectional groups
@@
Appendix D Text embeddings of adjectives sourced from two theories of social perception
Figure S.6: A: 3D visualization of text embeddings across theoretical models. Each marker denotes an adjective positioned by its PCA-reduced embedding. The center for each dimension, calculated as the mean of its items, is connected to its respective adjectives by dashed lines. B: Cosine similarities of text-to-text pairs. Within-valence similarities are indicated by the x-tick markers A+, A-, BP, BC, C+, C-. Cross-valence similarities are indicated by markers A, B, and C. Furthermore, the black dots indicate averages within dimension, the gray vertical segments indicate interquartile ranges and the gray horizontal line indicates the median of all dimensions. C: Heatmaps of Pearson correlation coefficients of positive and negative valence dimensions of the ABC model.
Table S.5: $p$-values from a one-sided t-test comparing within-valence cosine similarities (positive-negative) to cross-valence similarities. $p$-values $>$ 0.05 suggest that cross-valence similarities are not significantly smaller than within-valence, indicating the proximity of positive and negative valence items in the embedding space.
Category $+$/$P$-cross --/$C$-cross
Agency 0.29 0.80
Belief 0.23 0.59
Communion 0.56 0.31
This section investigates how the social perception dimensions are reflected within CLIP’s embedding space. To do this, we investigate within and cross-valence similarities. Specifically, let ${\bm{d}}^{+}_{1}$ and ${\bm{d}}^{+}_{2}$ be two text embeddings of adjectives [^3] In this subsection, we only use the adjectives as text prompt, without the sentence “A photo of a person” that was used to investigate text to image cosine similarities. of the Positive Agency dimension, and ${\bm{d}}^{-}_{1}$ and ${\bm{d}}^{-}_{2}$ be text embeddings of adjectives of the Negative Agency dimension. The within-valence similarity is given by calculating the cosine similarities $\cos{({\bm{d}}^{+}_{i},{\bm{d}}^{+}_{j})}$ for all positive adjective pairs $(i,j)$. The cross-valence similarity is determined by calculating $\cos{({\bm{d}}^{+}_{i},{\bm{d}}^{-}_{j})}$ for all mixed adjective pairs $(i,j)$.
Intuitively, we would assume that within-valence cosine similarities are higher than cross-valence similarities. However, $t$-tests comparing pairs of within- to cross-valence similarities indicate that cross-valence similarities are never significantly lower than their within-valence counterparts (Tab. S.5). In other words, opposing valences are close in space.
A 3D visualization of text embeddings (Fig. S.6A) supports these findings: opposing valences are closely clustered. In combination with CLIP’s modality gap between text and image embeddings [75], we can expect that cosine similarities between an image and a positive and negative adjective are positively correlated.
Furthermore, we investigate how these high correlations differ across dimensions and gender-race groups. To that end, we compute the Pearson correlation coefficients for each intersecting gender-race group across the three dimensions of Agency, Belief, and Communion. Fig. S.6C presents these coefficients in a heatmap. Contrary to Fig. S.6A, this figure includes image-text-cosine similarities, from which the neutral prompt is subtracted. Two key observations emerge from this analysis. First, within each heatmap, the highest magnitude is noted in the upper-mid, corresponding to White men. Second, across all heatmaps, Belief demonstrates the strongest correlation, whereas Agency exhibits the overall weakest correlation between positive and negative valence adjectives.
Unlike Agency and Communion, where descriptive adjectives often align along a positive-negative spectrum (such as friendly vs. unfriendly), Belief navigates ideological or value-based characteristics. This dimension is not framed in terms of positivity or negativity but rather in terms of conceptual oppositions, such as ‘progressive’ versus ‘conservative’. Fig. S.6C suggests a stronger correlation among these conceptual opposites within the Belief dimension compared to the positive and negative adjectives associated with Agency and Communion. It appears that when individuals assess traits related to Agency and Communion, there is a tendency to perceive these traits in more stark and definitive terms. On the other hand, Belief, with its ideological contrasts, seems to exhibit more of a gradient or continuum.
Appendix E Dataset visualizations
E.1 FairFace and UTKFace
@@
Out of the races present in FairFace, we used the original White and Black categories and aggregated East Asian and Southeastern Asian into a single category. Indian, Hispanic, and Middle Eastern were not used in this study. Similarly, for UTKFace, we use Asian, White, and Black but not the Indian and Others categories. For both datasets, we only used images annotated with age 20 or older. This is motivated by matching the general demographic distribution available in CausalFace.
Figure S.7: Example Images Left: FairFace, Right: UTKFace.
E.2 CausalFace
@@
The original CausalFace images have a resolution of 512x512 pixels. To emphasize the faces and minimize the influence of the background and clothing as well as to match the style of UTKFace images, we cropped each image to 432x432 pixels around the faces. This cropping was done by selecting pixels from 40 to 472 horizontally and from 0 to 432 vertically. An exception occurs with the side-view poses (Fig. S.10), where we horizontally cropped the images from 0 to 432 pixels for negative pose values and from 80 to 512 pixels for positive pose values, to ensure that faces are centered in the image.
@@
We also identified a slight spurious correlation between facial expression (“Smiling”) and the average image brightness in the original CausalFace images. To address this, we made small adjustments to the brightness of images in the Smiling category: First, we estimate the average brightness in the images’ face regions (binary masks for faces are available in CausalFace). Then, we scale the pixel values of “Smiling” variants to match the average brightness of the corresponding prototype image (where smiling=0, see Fig. S.9). We tested our results using both the adjusted and unadjusted images and observed only minor differences in CLIP cosine similarity when inferring with text prompts. Figs. S.8-S.11 show example images of the CausalFace images used in this study (after adjustments).
@@
Appendix F Brightness analysis in CausalFace
Motivated by the findings of Meister et al. [46], we explore the pixel-wise difference in grayscale values (brightness) between different demographic groups. More specifically, we examine whether the brightness of a pixel between two images (e.g., the male and female prototype within a seed) is greater, smaller, or equal. We do this for all seeds in the dataset and estimate the frequencies of how often these three cases occur. Formally, for each pixel position $P$, the displayed value $X_{P}$ that is shown in Figs. S.12 and S.13 is calculated as
@@ Appendix F Brightness analysis in CausalFace
Figure S.12: Pixel-wise differences in grayscale values between the male and female prototype within a seed aggregated across all seeds (see Eq. 12). Red means that these areas are more frequently darker for women, and blue areas are more often darker for men. As expected, the strongest effects are visible in the hair and beard regions. We attribute these regions to gender-specific attributes rather than unwanted confounds. We can also see a smaller effect for the background (slightly darker for women) and clothing (slightly darker for men). Even though the differences are small (±$\pm 4\%$ in brightness), future research could address eliminating these correlations.
@@
Appendix G Neutral text prompt associations in CausalFace
@@
Figure S.14: The neutral prompt bias. Probability densities for cosine similarities between text prompts and face images. Cosine similarities are shown separately for six demographic groups (colored lines). N=100 CausalFace face image prototypes (or pseudo-identities) are used for each demographic group (see also subsection E.1). The probability densities are approximated using kernel densities. Left column: Cosine similarities to a neutral prompt (e.g., “a photo of a person”). Notice that cosine similarities vary across prototype images (see also Fig. S.16). Center column: Cosine similarities to a prompt with a descriptive attribute (e.g., “a photo of a female person”). Right column: $\Delta$ cosine similarities after subtracting neutral prompt similarities for each image individually (see subsection 2.1), thus removing the bias for the neutral “person” association.

### v2

@@ A.1 Adapted SC-WEAT
Tab. S.1 shows the WEAT scores for CausalFace and the two observational datasets. The most striking similarities in means are observed when comparing Asians to Blacks and Whites to Blacks. In the comparison of Male versus Female, CausalFace and FairFace demonstrate closely matched scores, while UTKFace records a lower WEAT value.
A.2 Markedness
@@
A.3 Mean cosine similarities
@@
A.4 Skew@k, MaxSkew@k, NDKL
We present three additional bias metrics, following Geyik et al. [22].
@@ A.4 Skew@k, MaxSkew@k, NDKL
The leftmost block of the bottom subtable in Tab. S.1 depicts Skew@k. Asians predominantly receive positive values, indicative of a favorable social perception relative to other races. The values for Blacks, however, are distinctly negative. In the case of Whites, CausalFace consistently shows negative social perception, while FairFace and UTKFace mostly hover around zero. In contrast, Skew@k for men and women hovers near zero, implying no significant bias towards either gender.
@@
Appendix B Additional results
B.1 Facial expression, lighting, and pose confound social perception
@@
Fig. S.1 shows how facial expression, lighting, and pose affect social perception dimensions. A gray dashed line marks cosine similarity for a neutral prompt; deviations caused by valence attributes are shown as gray markers. Pearson correlations ($r$) support visual patterns (Tab. S.2).
@@
Since facial expressions are discussed in the main paper, we focus here on lighting and pose. The middle panel of Fig. S.1 illustrates the effect of lighting variations on social perception, revealing a consistent pattern across all dimensions. These variations, represented on the x-axis, correspond to different angles of light incidence, visualized in Fig. S.9. Although displayed on a linear scale for comparability, lighting variations are not inherently linear, so curve slopes are not central to our analysis. For all social dimensions, the highest scores are observed when the face is evenly lit (lighting = -1). Other configurations with high scores include lighting from above (lighting = 2 and 3). These findings are consistent with human subject studies [73, 74]. In contrast, the lowest scores are observed when the face is lit from below (lighting = 5 and 6), which tends to be perceived as unnatural by humans.
@@
The right panel of Fig. S.1 shows variations of $\Delta$ cosine similarity in pose. We observe a parallel trend in all social perception dimensions. The parallel trend is mirrored by a positive $r$ for all dimensions (average: rr=0.79). The curves are mirrored around a pose value of zero, representing a frontal stance. As a face tilts more to the left (pose << 0) or to the right (pose >> 0), the curves approach closer to the baseline. This effect is easier to interpret when considering that the raw (not the $\Delta$) cosine similarities instead. The raw values are the highest for frontal poses and decrease the more an image is tilted to the left or right (shaped like an upward-facing triangle). In other words, for side-pose photos, the image embedding is more distant from the “person” mode in the embedding space. This makes the interpretation of changes induced by adding social adjectives to the text prompt challenging. This observation suggests that uncontrolled pose variations in the observational, wild-collected datasets are a significant source of noise that makes analyzing other effects difficult and could explain a significant share of the noise present in wild-collected datasets (Appx. C). Future research could address this observation in more detail. Additionally, psychological literature indicates that head orientation (pose) can influence perceived emotions [75]. This would suggest that we should find a different impact on positive and negative attributes, at least for some of the varied head positions. However, in Zhang et al. [75], the postures affecting emotions were upward and downward head postures, while in CausalFace, we have left-right orientation variation, which may explain the difference in our observations. Additionally, we observe that negative valence is consistently ranked lower than positive valence, unaffected by head position. This suggests that another factor, currently unmeasured and uncontrolled, may affect this aspect.
@@
Figure S.1: Social perception as a function of Smiling, Lighting, and Pose. Curves for W (Warmth), C (Competence), A+, A- (Agency), BP, BC (Belief), C+, C- (Communion). Visualizations of the x-axis values are shown in Figs. S.7, S.8 and S.9. The cosine similarity between an image and a neutral text prompt is used as the reference and is depicted by the horizontal gray dashed line at 0. Markers are obtained by averaging across races and genders. Please see Figs. 3,5 for intersectional plots.
@@
B.2 Ageing influences social perception across intersectional groups
@@
Appendix C Dataset details
FairFace and UTKFace
@@ Appendix C Dataset details
Out of the races present in FairFace, we used the original White and Black categories and aggregated East Asian and Southeast Asian into a single category. Indian, Hispanic, and Middle Eastern were not used in this study. Similarly, for UTKFace, we use Asian, White, and Black, but not the Indian and Other categories. For both datasets, we only used images annotated with age 20 or older. This is motivated by matching the general demographic distribution available in CausalFace.
CausalFace
@@
The original CausalFace images have a resolution of 512x512 pixels. To emphasize the faces and minimize the influence of the background and clothing, as well as to match the style of UTKFace images, we cropped each image to 432x432 pixels around the faces. This cropping was done by selecting pixels from 40 to 472 horizontally and from 0 to 432 vertically. An exception occurs with the side-view poses (Fig. S.8), where we horizontally cropped the images from 0 to 432 pixels for negative pose values and from 80 to 512 pixels for positive pose values, to ensure that faces are centered in the image.
@@
We also identified a slight spurious correlation between facial expression (“Smiling”) and the average image brightness in the original CausalFace images. To address this, we made small adjustments to the brightness of images in the Smiling category: First, we estimate the average brightness in the images’ face regions (binary masks for faces are available in CausalFace). Then, we scale the pixel values of “Smiling” variants to match the average brightness of the corresponding prototype image (where smiling=0, see Fig. S.7). We tested our results using both the adjusted and unadjusted images and observed only minor differences in CLIP cosine similarity when inferring with text prompts. Figs. S.6-S.9 show example images of the CausalFace images used in this study (after adjustments).
@@
Appendix D Brightness analysis in CausalFace
Motivated by the findings of Meister et al. [48], we explore the pixel-wise difference in grayscale values (brightness) between different demographic groups. More specifically, we examine whether the brightness of a pixel between two images (e.g., the male and female prototype within a seed) is greater, smaller, or equal. We do this for all seeds in the dataset and estimate the frequencies of how often these three cases occur. Formally, for each pixel position $P$, the displayed value $X_{P}$ that is shown in Figs. S.10 and S.11 is calculated as
@@ Appendix D Brightness analysis in CausalFace
Figure S.10: Pixel-wise differences in grayscale values between the male and female prototype within a seed aggregated across all seeds (see Eq. 12). Red means that these areas are more frequently darker for women, and blue areas are more often darker for men. As expected, the strongest effects are visible in the hair and beard regions. We attribute these regions to gender-specific attributes rather than unwanted confounds. We can also see a smaller effect for the background (slightly darker for women) and clothing (slightly darker for men). Even though the differences are small ($\pm 4\%$ in brightness), future research could address eliminating these correlations.
@@
Appendix E Neutral text prompt associations in CausalFace
@@
Figure S.12: The neutral prompt bias. Probability densities for cosine similarities between text prompts and face images. Cosine similarities are shown separately for six demographic groups (colored lines). N=100 CausalFace face image prototypes (or pseudo-identities) are used for each demographic group (see also Appx. C). The probability densities are approximated using kernel densities. Left column: Cosine similarities to a neutral prompt (e.g., “a photo of a person”). Notice that cosine similarities vary across prototype images (see also Fig. S.14). Center column: Cosine similarities to a prompt with a descriptive attribute (e.g., “a photo of a female person”). Right column: $\Delta$ cosine similarities after subtracting neutral prompt similarities for each image individually (see subsection 2.1), thus removing the bias for the neutral “person” association.